OpenStudy (anonymous):

caleb kicked a ball. The function h(t)=-16t^2+80t describes the height of the ball, in feet, t seconds after he kicked it. How many seconds did the ball take to first reach 64 ft.

OpenStudy (regan):

solve 64=h(t)

OpenStudy (anonymous):

64 = -16t^2 + 80t 16t^2 - 80t + 64 = 0 16(t^2 - 5t + 4) = 0 16(t - 4)(t - 1) = 0 t = 1, 4 the first time is after 1 second